You are asked to evaluate the following two projects for the Norton corporation. Use a discount rate of 10 percent. Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods.

Project X (Videotapes
of the Weather Report)
($14,000 Investment)
Project Y (Slow-Motion
Replays of Commercials)
($34,000 Investment)
Year Cash Flow Year Cash Flow
1 $ 7,000 1 $ 17,000
2 5,000 2 10,000
3 6,000 3 11,000
4 5,600 4 13,000

  
a. Calculate the profitability index for project X. (Do not round intermediate calculations and round your answer to 2 decimal places.)

    

  
b. Calculate the profitability index for project Y. (Do not round intermediate calculations and round your answer to 2 decimal places.)


c. Which project would you select based on the profitability index?

  • Project X

  • Project Y
Answer #1

a)

Project X:

Profitability index = Present value / initial investment

Present value = 7,000 / (1 + 0.1)1 + 5,000 / (1 + 0.1)2 + 6,000 / (1 + 0.1)3 + 5,000 / (1 + 0.1)4

Present value = $18,418.8229

Profitability index = 18,418.8229 / 14,000

Profitability index = 1.32

b)

Project Y:

Profitability index = Present value / initial investment

Present value = 17,000 / (1 + 0.1)1 + 10,000 / (1 + 0.1)2 + 11,000 / (1 + 0.1)3 + 13,000 / (1 + 0.1)4

Present value = $40,862.6460

Profitability index = 40,862.6460 / 34,000

Profitability index = 1.20

c)

Project X as it has the higher profitability index